OpenStudy (anonymous):

You invest $2,200 at an interest rate that is compounded annually. After 8 years, your principal has increased to $3,128.62. What percent is the interest rate?

OpenStudy (anonymous):

much oblidged

OpenStudy (anonymous):

\[3128.62=2200\times (1+r)^8\] solve for r

OpenStudy (anonymous):

i would assume that u divide 2200 first which would give u 1.4221

OpenStudy (anonymous):

\[(1+r)^8=\frac{3128.62}{2200}\] \[1+r=\sqrt[8]{\frac{3128.62}{2200}}\]

OpenStudy (anonymous):

yeah divide. then take the 8th root. then subtract 1. then turn answer into a percent if that is what they want

OpenStudy (anonymous):

you will need a nice calculator to take the 8th root.

OpenStudy (anonymous):

root:8:(3128.62/2200)=1.0449999437

OpenStudy (anonymous):

so the percent would b 4%, right?

OpenStudy (anonymous):

i get the 8th root at 1.044999... so r = .044999 =4.45% rounded

OpenStudy (anonymous):

i would keep a couple decimal places, but yes, you have it

OpenStudy (anonymous):

it was 4.5 but at least i learned something

OpenStudy (anonymous):

excuse me i mean 4.5%

OpenStudy (anonymous):

i rounded incorrectly as you see. should round 4.4999 to 4.5